# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 5 Guidelines for Dataset Selection
Selecting an appropriate fairness evaluation dataset for LLMs requires aligning the choice with the model’s language scope, the evaluation task, desired control vs. realism, targeted bias types, and ethical considerations. This section offers evidence-based recommendations and a decision table to help researchers choose datasets—such as WinoBias, Winogender, GAP, BBQ, BOLD, StereoSet, CrowS-Pairs, and HolisticBias—suited to their evaluation scenarios. The recommendations build directly upon the bias types and estimation methods introduced in Sections 2 and 3.
@@
Language Scope (Monolingual vs. Multilingual). Most fairness datasets were originally developed in English. WinoBias, Winogender, GAP, BBQ, BOLD, StereoSet, and CrowS-Pairs are all English-based. For multilingual LLMs, recent adaptations such as French CrowS-Pairs or Korean BBQ (KoBBQ) may offer a starting point, but cultural translation and native-speaker validation are strongly advised. Researchers evaluating non-English models should ensure cultural validity by checking whether identity terms, stereotypes, or demographic groups are locally relevant.
Task Focus and Evaluation Format. Datasets should match the evaluation task format: (i) Coreference Resolution — Use WinoBias, Winogender, and GAP; (ii) Question Answering — Use BBQ; (iii) Stereotype Detection — Use CrowS-Pairs, StereoSet; (iv) Open-ended Generation — Use BOLD, HolisticBias.
Control vs. Ecological Validity. Researchers should decide whether their evaluation requires highly controlled conditions or realistic language usage scenarios. (i) Controlled Minimal Pairs: Datasets like WinoBias, Winogender, and CrowS-Pairs consist of synthetically constructed sentence pairs that differ only in a protected attribute (textite.g., “he” vs. “she”). This tight control ensures that any model output disparity can be directly attributed to the demographic variable, enabling precise statistical analysis. These datasets are ideal for isolating bias effects and performing significance tests with minimal confounding variables. (ii) Naturalistic Prompts: Datasets such as GAP, BOLD, and parts of BBQ use real-world or plausibly written prompts. This format better reflects real user interactions and thus provides greater ecological validity. For instance, GAP offers authentic Wikipedia sentences with naturally occurring ambiguous pronouns, while BOLD prompts derive from encyclopedic contexts across five social domains. (iii) Balanced Approaches: Some datasets—especially BBQ—strive for a hybrid structure: prompts are constructed but contextually realistic, and bias is probed both under ambiguous and disambiguated conditions. This makes such datasets valuable for evaluations that require both control and realism.
In practice, combining both types of datasets provides complementary evidence. Controlled datasets allow targeted diagnosis of bias, while naturalistic datasets test whether those biases emerge under real-world usage. The combination enhances both validity and generalizability.
@@
Bias Dimensions and Coverage. Fairness evaluation datasets differ in the demographic dimensions they target. Researchers should choose datasets whose protected attributes align with the specific bias categories under investigation. (i) Single-Dimension Datasets (textite.g., Gender): Datasets like WinoBias, Winogender, and GAP focus exclusively on gender. WinoBias and Winogender test whether coreference or pronoun resolution systems behave differently based on gendered pronouns in synthetically constructed sentences. GAP offers naturally occurring sentences with ambiguous gendered pronouns, supporting fairness evaluations grounded in real-world data. These datasets are ideal for studies that target binary gender bias in isolation but do not generalize to other bias axes. (ii) Multi-Dimensional Datasets: When the goal is to audit biases across multiple protected attributes (textite.g., race, religion, socioeconomic status), more comprehensive datasets are preferable. BBQ and CrowS-Pairs each cover nine social dimensions, including race, gender, religion, nationality, age, sexual orientation, disability, physical appearance, and socioeconomic background. These datasets are suitable for evaluating intersectional and structural model biases, particularly when multiple demographic categories interact in the same input. (iii) Fine-Grained and Inclusive Coverage: HolisticBias extends beyond conventional categories, covering 13 demographic axes and nearly 600 identity descriptors. These include underrepresented dimensions such as neurodiversity, body type, cultural background, and nonbinary gender identities. The dataset is designed to surface subtle representational imbalances and is particularly useful when researchers aim to audit models for inclusivity and coverage breadth. Since it is built from templates, HolisticBias can also be filtered to focus on specific subgroups or identity intersections. (iv) Dataset Selection Based on Target Bias Type: Researchers concerned with a specific bias—such as occupational gender stereotypes, racial polarity in sentiment, or nationality-linked toxicity—should consult the dataset’s documented taxonomy and demographic labels. For instance, to probe profession-related gender bias, StereoSet and BOLD are well-suited. For testing national origin or racial stereotypes, CrowS-Pairs and BBQ provide explicit group comparisons in stereotype contexts. For broad inclusivity studies, HolisticBias offers the highest demographic resolution.
@@
In summary, dataset selection should be driven by the targeted fairness axis. Using a gender-only benchmark may provide misleading confidence if other demographic biases remain untested. Comprehensive audits benefit from datasets with wide demographic coverage and interpretable group labels.
@@
Ethical and Practical Considerations. Beyond technical fitness, dataset selection for fairness evaluation must consider ethical risks, annotator quality, and resource constraints. These factors can significantly influence the reliability and responsibility of fairness assessments. (i) Cultural Validity and Transferability: Many datasets—such as WinoBias, CrowS-Pairs, and BBQ—were constructed based on U.S.-centric stereotypes and demographic categories. When evaluating LLMs in non-Western or multilingual contexts, researchers should examine whether these datasets reflect locally meaningful groupings, language use, and social dynamics. For cross-lingual applications, datasets like KoBBQ or French CrowS-Pairs offer culturally adapted variants. Where no localized dataset exists, translation and adaptation should be guided by native speakers to ensure relevance and fairness. (ii) Annotator and Label Quality: Crowdsourced datasets vary in annotation reliability. For example, CrowS-Pairs was constructed with strict validation filters and achieved higher agreement rates than StereoSet, which exhibited inconsistencies due to more permissive quality control. When precision is important—textite.g., in minimal-pair evaluation—datasets with rigorous annotation pipelines should be preferred. For generated labels (textite.g., toxicity scores), researchers should audit the metric itself (as described in Section 2.3) to ensure the instrument does not introduce bias. (iii) Exposure to Harmful Content: Some datasets deliberately include offensive, stereotypical, or discriminatory language to test model sensitivity. This is common in CrowS-Pairs, StereoSet, and some examples in BBQ. While necessary for stress testing, such content carries ethical risks, including potential harm to annotators, readers, or model end-users. Researchers should apply appropriate safeguards: clearly flag harmful content, avoid using such data during training, and ensure outputs are not unintentionally redistributed without context. (iv) Model Access and Metric Compatibility: Certain datasets require internal model access (textite.g., token-level likelihoods) to compute bias scores—StereoSet, for instance, uses a scoring scheme based on log-likelihood differences. This poses challenges for closed models (textite.g., GPT-4 via API). In contrast, datasets like BBQ and HolisticBias can be evaluated with only the generated output, using external metrics such as sentiment, regard, or toxicity. Dataset–metric compatibility must be considered when model access is limited. (v) Computational Budget and Dataset Scale: Datasets vary greatly in size. HolisticBias includes over 450k prompts, offering wide coverage but requiring significant inference time and storage. Others, such as Winogender (720 examples) or CrowS-Pairs (1.5k pairs), are compact and suited for rapid or exploratory testing. Researchers with limited computational budgets should select smaller datasets or sample larger corpora proportionally to bias category coverage.
@@
Multilingual bias evaluation Adapted CrowS-Pairs (textite.g., French), KoBBQ, translated HolisticBias
@@
Table 9 offers a high-level summary of these recommendations. This selection framework serves as a bridge between dataset properties (analyzed in Sections 3–5) and practical deployment in fairness evaluation pipelines.
@@ 6 Conclusion
This survey provides the first systematic investigation into the datasets used to evaluate fairness in large language models, shifting the focus from model behaviour to the benchmarks themselves. By proposing a unified taxonomy and a general experimental methodology, we enable consistent, reproducible analysis of dataset-level biases across both counterfactual and prompt-based corpora. Our findings show that many widely used benchmarks—despite being designed to test bias—exhibit structural and statistical imbalances, including representativeness gaps, annotation artifacts, stereotype leakage, and differential metric bias. These issues, if unexamined, can distort fairness claims and limit the generalizability of results. Through empirical audits of diverse datasets, we demonstrate that fairness evaluation must begin with fair, well-understood benchmarks. We encourage the community to routinely assess dataset integrity alongside model outputs, validate the tools used for scoring bias, and adopt diverse, multilingual, and context-sensitive datasets. By treating fairness datasets as first-class research objects, we take a critical step toward more robust, inclusive, and accountable language model evaluation.

### v2

@@ 5 Guidelines for Dataset Selection
Selecting an appropriate fairness evaluation dataset for LMs requires aligning the choice with the model’s language scope, the evaluation task, desired control vs. realism, targeted bias types, and ethical considerations. This section offers evidence-based recommendations and a decision table to help researchers choose datasets suited to their evaluation scenarios. The recommendations build directly upon the bias types and estimation methods introduced in Sections 3 and 4.
@@
Language Scope (Monolingual vs. Multilingual). Most fairness datasets were originally developed in English. WinoBias, Winogender, GAP, BBQ, BOLD, StereoSet, and CrowS-Pairs are all English-based. For multilingual LMs, recent adaptations such as French CrowS-Pairs or Korean BBQ (KoBBQ) may offer a starting point, but cultural translation and native-speaker validation are strongly advised. Researchers evaluating non-English models should ensure cultural validity by checking whether identity terms, stereotypes, or demographic groups are locally relevant.
Task Focus and Evaluation Format. Researchers should align the dataset with the evaluation task that best captures the behaviour they wish to measure. For coreference and pronoun resolution, the syntactically controlled WinoBias, WinoGender, and GAP datasets remain the most precise instruments for detecting gender–occupation asymmetries in antecedent assignment. When the objective is to probe stereotypical reasoning in multiple‑choice question answering, BBQ and UnQover supply carefully balanced option sets that reveal whether a model preferentially selects stereotype‑consistent answers. The study of stereotype association or classification biases is best served by likelihood‑ or accuracy‑based resources such as CrowS‑Pairs, StereoSet, RedditBias, and the template‑rich HolisticBias, each enabling quantification of demographic disparities under minimal textual changes. Finally, the evaluation of open‑ended generation requires corpora that elicit free text; BOLD, RealToxicityPrompts, HONEST, and TrustGPT expose emergent toxicity, regard differentials, and stereotype leakage that are invisible in constrained settings. Selecting a dataset whose structure mirrors the intended task therefore maximises construct validity and interpretability of any fairness metric.
Control vs. Ecological Validity. A fundamental design choice concerns the trade‑off between experimental control and linguistic realism. Controlled minimal‑pair corpora, exemplified by WinoBias, CrowS‑Pairs, HolisticBias, and the counterfactual slices of UnQover, vary only a single protected attribute within otherwise identical contexts. Such tight control limits confounding factors, so performance gaps can be attributed with high confidence to the demographic variable, enabling rigorous statistical testing. In contrast, naturalistic prompts such as those found in GAP, BOLD, RealToxicityPrompts, and RedditBias embed demographic information in authentic discourse. These datasets confer ecological validity, revealing whether biases diagnosed under laboratory conditions surface in real‑world usage and uncovering compound harms, like toxic language amplification, that synthetic pairs cannot capture. Between these poles lie hybrid resources like BBQ and HONEST, whose handcrafted yet contextually rich prompts retain demographic balance while preserving linguistic variability; they reconcile the precision of templates with the expressive breadth of natural language.
Combining at least one controlled dataset with one naturalistic or hybrid corpus provides complementary evidence: the former isolates causal bias mechanisms, whereas the latter tests the robustness and generalisability of mitigation strategies under realistic deployment conditions. A balanced evaluation suite therefore strengthens both internal and external validity in fairness research.
@@
Bias Dimensions and Coverage. Fairness evaluation datasets differ in the demographic dimensions they target. Researchers should choose datasets whose protected attributes align with the specific bias categories under investigation. (i) Single-Dimension Datasets (e.g., Gender): Datasets like WinoBias, Winogender, and GAP focus exclusively on gender. WinoBias and Winogender test whether coreference or pronoun resolution systems behave differently based on gendered pronouns in synthetically constructed sentences. GAP offers naturally occurring sentences with ambiguous gendered pronouns, supporting fairness evaluations grounded in real-world data. These datasets are ideal for studies that target binary gender bias in isolation but do not generalize to other bias axes. (ii) Multi-Dimensional Datasets: When the goal is to audit biases across multiple protected attributes (e.g., race, religion, socioeconomic status), more comprehensive datasets are preferable. BBQ and CrowS-Pairs each cover nine social dimensions, including race, gender, religion, nationality, age, sexual orientation, disability, physical appearance, and socioeconomic background. These datasets are suitable for evaluating intersectional and structural model biases, particularly when multiple demographic categories interact in the same input. (iii) Fine-Grained and Inclusive Coverage: HolisticBias extends beyond conventional categories, covering 13 demographic axes and nearly 600 identity descriptors. These include underrepresented dimensions such as neurodiversity, body type, cultural background, and nonbinary gender identities. The dataset is designed to surface subtle representational imbalances and is particularly useful when researchers aim to audit models for inclusivity and coverage breadth. Since it is built from templates, HolisticBias can also be filtered to focus on specific subgroups or identity intersections. (iv) Dataset Selection Based on Target Bias Type: Researchers concerned with a specific bias—such as occupational gender stereotypes, racial polarity in sentiment, or nationality-linked toxicity—should consult the dataset’s documented taxonomy and demographic labels. For instance, to probe profession-related gender bias, StereoSet and BOLD are well-suited. For testing national origin or racial stereotypes, CrowS-Pairs and BBQ provide explicit group comparisons in stereotype contexts. For broad inclusivity studies, HolisticBias offers the highest demographic resolution.
@@
In summary, dataset selection should be driven by the targeted fairness axis. Using a gender-only dataset may provide misleading confidence if other demographic biases remain untested. Comprehensive audits benefit from datasets with wide demographic coverage and interpretable group labels.
@@
Ethical and Practical Considerations. Beyond technical fitness, dataset selection for fairness evaluation must consider ethical risks, annotator quality, and resource constraints. These factors can significantly influence the reliability and responsibility of fairness assessments. (i) Cultural Validity and Transferability: Many datasets—such as WinoBias, CrowS-Pairs, and BBQ—were constructed based on U.S.-centric stereotypes and demographic categories. When evaluating LLMs in non-Western or multilingual contexts, researchers should examine whether these datasets reflect locally meaningful groupings, language use, and social dynamics. For cross-lingual applications, datasets like KoBBQ or French CrowS-Pairs offer culturally adapted variants. Where no localized dataset exists, translation and adaptation should be guided by native speakers to ensure relevance and fairness. (ii) Annotator and Label Quality: Crowdsourced datasets vary in annotation reliability. For example, CrowS-Pairs was constructed with strict validation filters and achieved higher agreement rates than StereoSet, which exhibited inconsistencies due to more permissive quality control. When precision is important—e.g., in minimal-pair evaluation—datasets with rigorous annotation pipelines should be preferred. For generated labels (e.g., toxicity scores), researchers should audit the metric itself (as described in Section 2.3) to ensure the instrument does not introduce bias. (iii) Exposure to Harmful Content: Some datasets deliberately include offensive, stereotypical, or discriminatory language to test model sensitivity. This is common in CrowS-Pairs, StereoSet, and some examples in BBQ. While necessary for stress testing, such content carries ethical risks, including potential harm to annotators, readers, or model end-users. Researchers should apply appropriate safeguards: clearly flag harmful content, avoid using such data during training, and ensure outputs are not unintentionally redistributed without context. (iv) Model Access and Metric Compatibility: Certain datasets require internal model access (e.g., token-level likelihoods) to compute bias scores—StereoSet, for instance, uses a scoring scheme based on log-likelihood differences. This poses challenges for closed models (e.g., GPT-4 via API). In contrast, datasets like BBQ and HolisticBias can be evaluated with only the generated output, using external metrics such as sentiment, regard, or toxicity. Dataset–metric compatibility must be considered when model access is limited. (v) Computational Budget and Dataset Scale: Datasets vary greatly in size. HolisticBias includes over 450k prompts, offering wide coverage but requiring significant inference time and storage. Others, such as Winogender (720 examples) or CrowS-Pairs (1.5k pairs), are compact and suited for rapid or exploratory testing. Researchers with limited computational budgets should select smaller datasets or sample larger corpora proportionally to bias category coverage.
@@
Multilingual bias evaluation Adapted CrowS-Pairs (e.g., French), KoBBQ, translated HolisticBias
@@
Table 9 offers a high-level summary of these recommendations. This selection framework serves as a bridge between dataset properties (analyzed in Sections 3) and  4 and practical deployment in fairness evaluation pipelines.
6 Future Challenges
Despite consolidating existing datasets and proposing a unified lens for dataset-level bias, substantial challenges remain. First, the community needs fairness datasets with broader linguistic and cultural coverage. Most existing resources are English-centric and Western-oriented, which constrains external validity. Future datasets should be multilingual and culturally grounded, capturing grammatical gender systems, rich morphology, dialectal variation, and code-switching phenomena, with particular attention to low-resource languages. Doing so would enable cross-lingual comparability and help surface fairness failures that only emerge under specific linguistic structures or sociocultural norms.
A second challenge concerns intersectionality and multi-attribute fairness. Many current datasets vary only one protected attribute at a time, which obscures compounding effects (e.g., gender and race) that are common in real deployments. Next-generation resources should systematically vary multiple attributes jointly while balancing two desiderata: (i) reflecting realistic co-occurrence patterns in target populations (to avoid misleading priors) and (ii) preserving counterfactual control (to enable clean causal probes). Meeting this goal will likely require participatory design, careful privacy protections, and explicit documentation of sampling rationales and limitations.
Sustained governance is equally important. Social norms, demographic baselines, and model capabilities evolve, so fairness datasets must evolve with them. We see a need for community-maintained governance frameworks that cover versioning, transparent documentation, auditable change logs, and clear procedures to mitigate dataset overfitting. Open contribution pathways (e.g., participatory annotation, challenge sets, issue tracking) and reproducible evaluation pipelines can help keep datasets relevant and trustworthy over time.
Finally, closing the gap between dataset scores and real-world impact remains a pressing priority. Future work should link dataset-level fairness signals to domain-specific harms and benefits, especially in high-stakes settings such as healthcare, education, and law. This entails designing evaluation protocols that are sensitive to contextual risk, articulating decision thresholds that translate into operational safeguards, and aligning measurement with governance artifacts used in practice (compliance checks, procurement requirements, incident response). We hope these challenges will catalyze collaborative efforts toward the next generation of fairness datasets and evaluation practices that are both technically rigorous and socially responsive.
@@ 7 Conclusion
This survey provides the first comprehensive examination of the datasets used to assess fairness in language models, shifting the analytical focus from model outcomes to the data that underwrite them. We propose a unified taxonomy that distinguishes constrained-form and open-ended evaluation settings, and we formalize an experimental protocol that operationalizes dataset-level biases, including representativeness, annotation bias, and stereotype leakage, within a common statistical framework. Applying this framework, our empirical audits show that several widely used datasets, even those explicitly designed to probe bias, embed structural and statistical imbalances that can distort fairness conclusions and limit external validity. These findings indicate that rigorous fairness research must begin with well-characterized, methodologically sound datasets; in practice this calls for routine dataset audits, careful validation of scoring instruments, and sustained investment in multilingual, intersectional, and context-sensitive resources. Building on this foundation, the next section outlines the future challenges that must be addressed to achieve robust, inclusive, and accountable evaluation, with particular attention to expanding linguistic and cultural coverage, modeling intersectional identities, establishing durable governance and maintenance practices, and linking dataset signals to real-world impact.